OpenStudy (anonymous):

find the real numbers x and y that make the equation true. 3+yi = x-7i

OpenStudy (ghazi):

3+yi=x-7i x-yi=3+7i compare real and imaginary part.... x=3 and -y=7 or y=-7

OpenStudy (anonymous):

oh ok so set it up like the same. ok gotcha. thanks dude

OpenStudy (ghazi):

:) YW